OpenStudy (yttrium):

The 60% of 120% of 45% of 81% of the first number is 27% of 100% of 54% of 70% of the second number. What are the two numbers? I thought it has something to do with split-difference though I know nothing about it.

OpenStudy (atlas):

60% means 0.6 and so on so 0.6 *1.2*0.45*0.81 of a number = 0.27*1*0.54*0.7 of another number

OpenStudy (atlas):

2nd number/1st number = (0.6*1.2*0.45*0.81)/(0.27*1*0.54*0.7)